# Laura Barros        Aiko Schmeißer

### v3

@@ 2.2 Local councilors
As in many other contexts, serving as a local councilor in Brazil can be financially rewarding. [^7] According to PoliticalSalaries.com, a not-for-profit open data project, being a legislator is financially attractive in many countries, as measured by legislator salaries as a percentage of GDP per capita. In Brazil, legislators earn salaries equivalent to 444% of GDP per capita, a ratio similar to other upper-middle-income countries, such as Colombia (490%) and South Africa (443%). The wages of councilors are substantial on average – amounting to about 2.5 times the average wage in their municipalities. Councilor wages also vary strongly by municipality size – being capped at 20% of state deputies’ wages in municipalities with less than 10,000 inhabitants and at 75% in municipalities with more than 500,000 inhabitants. Importantly, serving as a councilor is not a full-time activity and allows legislators to retain their private-sector jobs after being elected. On average, legislators need to be present in the council only four days per month (Ferraz and Finan, 2011). Besides the salaries received by elected councilors, patronage channels can also generate returns for candidates who are not elected. Colonnelli et al. (2020) find that non-elected councilor candidates are about 15 percentage points more likely to be employed in the public sector and receive 40% higher earnings in the formal economy if their supported party comes into power.
@@
How large are the financial costs associated with candidacies for local council? To finance their campaigns, candidates rely on party funds, financial contributions from individuals and firms, and their own resources. [^10] Campaign spending limits were introduced in 2015 (Avis et al., 2022). In addition to banning financial contributions from private firms, the electoral law also set spending limits for political campaigns based on the population of the municipality. Since 2019, candidates for local council are only allowed to fund up to 10% of their municipality’s spending limit with their own resources, ensuring more equitable running conditions. We collected data from the Superior Electoral Court (Tribunal Superior Eleitoral, TSE) on all campaign contributions received by councilor candidates in the electoral cycles 2004 to 2020. Appendix Table A.1 presents detailed summary statistics for each year. Across all years, the median candidate raised a total of 1,587 BRL (392 USD), with only 168 BRL (34 USD) coming from own resources. Only about half of all candidates use any private funds for the campaign. These figures suggest that personal financial contributions play a limited role in local campaigns, suggesting that the financial barriers to running for councilor are relatively low.
@@ 2.3 Theoretical considerations
We now introduce a simple economic framework to illustrate how a layoff may affect an individual’s likelihood of becoming a party member or councilor candidate. [^11] For economic models of candidate selection, see, for example, Black (1972), Caselli and Morelli (2004), and Messner and Polborn (2004). Consider an individual who decides how to divide her non-leisure time between pursuing a career in the private sector or in politics. Let $h^{P}$ denote the share of time devoted to politics. If the individual does not engage in politics ($h^{P}=0$), she earns a market hourly wage of $w^{M}$. When entering politics ($0<h^{P}\leq 1$), she foregoes a share of her earnings that is proportional to $h^{P}$. [^12] Since neither party membership nor local councilor is a full-time political commitment, we consider the possibility of simultaneously working in the private sector, albeit with a reduced number of hours. At the same time, political engagement provides her with an array of expected benefits, $b^{P}$, which increase with the time devoted to political activities. These benefits include monetary rewards, such as councilor salaries when elected to office (Ferraz and Finan, 2011), public sector employment opportunities resulting from political connections to the winning party or candidate (Brollo et al., 2017; Colonnelli et al., 2020; Barbosa and Ferreira, 2023), and improved employment prospects in the private labor market under co-partisan employers (Colonnelli et al., 2022). Non-monetary rewards, such as ego rents or prosocial benefits, may also play a role (Gulzar and Khan, 2024). Finally, to engage in politics, the individual incurs entry costs $c^{P}$ (e.g., campaign costs or membership registration fees).
@@
Second, the effects increase with the expected benefits of political engagement during unemployment, $(1-\lambda)b^{P}$, weighted by the time not already allocated to politics, $1-h^{P}$. Intuitively, this term reflects the notion that, when remaining unemployed, individuals can reallocate time that would otherwise be spent in market work towards political activities, thereby gaining additional political benefits. To shed light on one of the monetary returns from engaging with politics, we will investigate whether the treatment effects are larger in localities where councilors earn higher salaries. [^14] In addition to monetary returns, political involvement may also provide policy payoffs. In citizen-candidate models (Osborne and Slivinski, 1996; Besley and Coate, 1997), political candidates represent particular policies and care about winning per se. As argued by Hall (2019) and Thomsen (2017), the policy returns may be especially relevant for ideologically extreme candidates. Thus, we will examine potential heterogeneity in the effects on memberships and candidacies in parties of different ideological positions. Moreover, we will test the role of time resources by exploiting discontinuities in eligibility for UI benefits that increase laid-off workers’ incentives to remain unemployed.
@@
The framework also provides predictions about the potential implications of layoffs for the quality of individuals who select into politics. Given that the participation effects of layoffs are predicted to be larger for individuals with higher income losses and higher expected benefits from political engagement, the consequences for selection depend on how these factors are related to individual competence. High-ability individuals – such as those with higher education – tend to earn higher wages in the private sector. Thus, a layoff may represent a larger shock to their opportunity costs, leading to a more pronounced increase in their incentives to enter politics. In addition, the expected benefits of political engagement may vary with individuals’ competence. For instance, the returns to running for office rise with the probability of winning, which may be higher for more capable individuals. Moreover, the expected likelihood of gaining access to a public-sector job through political connections may depend on ability, either negatively if political support acts as a substitute for individual competence (Colonnelli et al., 2020) or positively if politicians use their discretionary power to appoint more competent individuals (Brollo et al., 2017; Bazzi and Labanca, 2023).
@@ 3.2 Labor market data
We obtain matched employer-employee data covering the universe of formal employees between 2004 and 2018 from the Relação Anual de Informações Sociais (RAIS), an administrative database collected annually by the Ministry of Labor and Employment. [^15] In 2015, formally employed individuals accounted for approximately 55% of the Brazilian labor force (Derenoncourt et al., 2021). The database was created in 1975 for statistical and administrative purposes and contains detailed information such as hiring and separation date, average monthly earnings, contracted hours, occupation, firm’s location and industry, and worker’s socio-demographic characteristics. For our analysis, we make use of the identified dataset, containing individuals’ names, taxpayer registry numbers (CPF), as well as firms’ tax identification numbers (Cadastro Nacional da Pessoa Jurídica, CNPJ). This enables us to identify individuals’ employment histories over time and across firms.
@@ 3.3 Merging political and labor market data
We merge the information on political candidates with the employment data using an individual’s taxpayer registry number (CPF). As for party membership, we merge this information to the RAIS data based on each individual’s full name. To minimize measurement error when exactly matching on names, we follow Britto et al. (2022a, b) and Bhalotra et al. (2025) and restrict our sample to individuals who have a unique name in Brazil. Individuals with unique names cover roughly half of the Brazilian population (50.2% in our sample), given that individuals commonly have several surnames, including at least one from both their father and mother. [^16] To identify unique names, we build a dataset of names in Brazil. We collect the names of all workers who appear in the RAIS data in any year between 2004 and 2018, and supplement this list with information on the universe of beneficiaries in the social programs Bolsa Família and Benefício de Prestação Continuada between 2013 and 2021 and on all party members between 2000 and 2020. The resulting list covers about 90% of the Brazilian adult population. Not observing the names of a small part of the population will introduce some measurement error in our merged outcome variables which is, however, not expected to be related to the likelihood of being laid-off.
@@ 4.1 Empirical strategy
The main difficulty in estimating the effect of layoffs on individuals’ entry into politics is the fact that layoffs are not random. On the one hand, individual shocks could simultaneously influence both political participation and the likelihood of losing a job, leading to a spurious relationship. On the other hand, individuals’ political behavior could also affect their probability of being dismissed. Colonnelli et al. (2022) show, for instance, that political discrimination affects labor market outcomes in Brazil. To address these identification challenges, we exploit the timing of firm-level mass layoffs as a source of plausibly exogenous variation in individual-level job loss using a difference-in-difference design in a matched sample of workers. The timing of mass layoffs is arguably orthogonal to individual workers’ political behavior and other confounding worker-level shocks, and has been widely used to estimate the effects of job loss on various outcomes. In particular, our empirical strategy closely follows recent papers that study the effect of dismissals in Brazil (Bhalotra et al., 2025; Britto et al., 2022b, a; Amorim et al., 2025). This subsection describes how we construct the estimation sample and then presents the estimation model.
@@
Sample selection. We restrict our sample to workers aged 25 to 50 with a full-time (i.e., at least 30 hours per week), open-ended, private-sector job of at least six months tenure. Among those, the treatment group comprises all workers displaced without just cause from a mass layoff firm between 2009 and 2011.171717We exclude individuals who are recalled by their initial employer within five years after the layoff. In cases where individuals experienced multiple mass layoffs between 2009 and 2011, we keep only the first one. This treatment period allows us to study long-run effects over three election cycles (12 years) after the layoff, and also to check for pre-trends over two election cycles (8 years) before the layoff. In our baseline specification, we define a mass layoff firm as a firm with at least 30 workers dismissing at least 30% of its workforce in a given year, and later we show that our results are robust to various thresholds of firm size and layoff share. [^18] We always drop firms from the mass layoff sample if they reallocate under a new identifier, specifically if at least 30% of their dismissed workers transition to the same new firm identifier in the following year.
@@
We then match each treated individual with a control worker who is not dismissed in the same year and who works in a non-mass layoff firm. [^19] Note that even though control workers are not laid off in the matching year, they may be laid off in later years. Given that we estimate effects for up to 12 years after the layoff, restricting control workers to those who were continuously employed throughout the whole observation period would result in a small and highly selected pool of control workers. Defining control workers without conditioning on future employment follows a large share of papers in the job loss literature (e.g., Britto et al., 2022b; Bertheau et al., 2023; Schmieder et al., 2023). In addition, we do not impose any restrictions on potential layoffs prior to the matching year, other than the minimum six-month tenure requirement. Given the high turnover rates in Brazil, stricter tenure requirements would yield a substantially selected sample. In Appendix Figure B.1, we show that treated and matched control individuals were not differently exposed to mass layoffs before the treatment period. To obtain a comparable control group, we perform an exact matching on gender, education (5 categories), age (deciles), monthly earnings (deciles), tenure (deciles), firm size (deciles), and state (27 categories). [^20] We allow control individuals to be matched to more than one treated individual. In cases where more than one control individual was matched to one treated individual, we randomly selected one control individual. For 79.5% of the workers in the initial treatment pool, we are able to find a comparable control worker. We end up with an estimation sample that consists of nearly one million treated and matched control workers.
@@
Table 1 presents summary statistics for treated and control individuals and shows that both groups are similar in socio-demographic, job, and firm characteristics. For all considered variables, the standardized mean difference between the two groups is well below the threshold of 0.20 recommended by Imbens and Rubin (2015). Importantly, treated and control units are also balanced in terms of pre-treatment political outcome variables which were not part of the matching process. In both groups, approximately 7% of individuals were members of a political party in 2008. Nearly 2% newly registered as a member in the election cycle between 2005 and 2008, and 0.1% ran for local councilor in the 2008 election. [^21] Appendix Table B.1 also shows that there are no differences in the ideology, local incumbency, and union affiliation of the parties that treated and control observations affiliated with or ran for before the layoff. The validity of our DiD design does not require that treated and control individuals are similar in levels, but covariate balance makes it more plausible that both groups exhibit similar trends in counterfactual outcomes.
@@
which summarizes the average treatment effects over all periods in the coefficient $\delta_{DiD}$. In both models, we include municipality $\times$ two-digit industry $\times$ cycle fixed effects, $\mu_{m(i)k(i)c}$, which flexibly control for local and industry shocks that may simultaneously cause mass layoffs and changes in political behavior. Importantly, the municipality $\times$ cycle fixed effects also hold constant the local political environment and allow us to separate individual-level supply-side effects from potential effects of mass layoffs on the demand for politicians among the local electorate. [^22] Note that the estimation models (1) and (2) rely on a single treatment period, as all treated workers were laid off in the 2012 cycle. Thus, our main specification is not susceptible to weighting issues that arise in staggered difference-in-differences models with heterogeneous treatment effects (de Chaisemartin and D’Haultfœuille, 2020; Goodman-Bacon, 2021; Sun and Abraham, 2021).
@@
where $i$, $t$, and $t^{*}$ refer to workers, calendar years, and layoff years, respectively. [^24] For control workers, we assign the layoff year of their matched treated worker. $\delta_{k}$ capture the dynamic treatment effects for each event year after the layoff $k$. The model includes calendar year $\times$ layoff year fixed effects, $\lambda_{tt^{*}}$, to ensure that the identifying variation only comes from comparing treated to control individuals who have been matched in the same year. [^25] This stacking approach avoids the “forbidden comparisons” between late and early-treated individuals that can cause negative weights under heterogeneous treatment effects (Cengiz et al., 2019; Dube et al., 2023).

### v4

@@ 2.2 Local councilors
As in many other contexts, serving as a local councilor in Brazil can be financially rewarding. [^8] According to PoliticalSalaries.com, a not-for-profit open data project, being a legislator is financially attractive in many countries, as measured by legislator salaries as a percentage of GDP per capita. In Brazil, legislators earn salaries equivalent to 444% of GDP per capita, a ratio similar to other upper-middle-income countries, such as Colombia (490%) and South Africa (443%). The wages of councilors are substantial on average – amounting to about 2.5 times the average wage in their municipalities. Councilor wages vary strongly by municipality size – being capped at 20% of state deputies’ wages in municipalities with less than 10,000 inhabitants and at 75% in municipalities with more than 500,000 inhabitants. Importantly, serving as a councilor is not a full-time activity and allows legislators to retain their private-sector jobs after being elected. On average, legislators need to be present in the council only four days per month (Ferraz2011). Besides the salaries received by elected councilors, patronage channels can also generate returns for candidates who are not elected. colonnelli2020 find that non-elected councilor candidates are about 15 percentage points more likely to be employed in the public sector and receive 40% higher earnings in the formal economy if their supported party comes into power.
@@
How large are the financial costs associated with candidacies for local council? To finance their campaigns, candidates rely on party funds, financial contributions from individuals and firms, and their own resources. [^11] Campaign spending limits were introduced in 2015 (avis2022). In addition to banning financial contributions from private firms, the electoral law also set spending limits for political campaigns based on the population of the municipality. Since 2019, candidates for local council are only allowed to fund up to 10% of their municipality’s spending limit with their own resources, ensuring more equitable running conditions. We collected data from the Superior Electoral Court (Tribunal Superior Eleitoral, TSE) on all campaign contributions received by councilor candidates in the electoral cycles 2004 to 2020. Appendix Table A.1 presents detailed summary statistics for each year. Across all years, the median candidate raised a total of 1,587 BRL (392 USD), with only 168 BRL (34 USD) coming from own resources. Only about half of all candidates use any private funds for the campaign. These figures suggest that personal financial contributions play a limited role in local campaigns, suggesting that the financial barriers to running for councilor are relatively low.
@@ 2.3 Theoretical considerations
We now introduce a simple economic framework to illustrate how a layoff may affect an individual’s likelihood of becoming a party member or councilor candidate. [^12] For economic models of candidate selection, see, for example, Black1972, Caselli2004, and Messner2004. Consider an individual who decides how to divide her non-leisure time between pursuing a career in the private sector or in politics. Let $h^{P}$ denote the share of time devoted to politics. If the individual does not engage in politics ($h^{P}=0$), she earns a market hourly wage of $w^{M}$. When entering politics ($0<h^{P}\leq 1$), she foregoes a share of her earnings that is proportional to $h^{P}$. [^13] Since neither party membership nor local councilor is a full-time political commitment, we consider the possibility of simultaneously working in the private sector, albeit with a reduced number of hours. At the same time, political engagement provides her with an array of expected benefits, $b^{P}$, which increase with the time devoted to political activities. These benefits include monetary rewards, such as councilor salaries when elected to office (Ferraz2011), public sector employment opportunities resulting from political connections to the winning party or candidate (brollo2017; colonnelli2020; barbosa2023), and improved employment prospects in the private labor market under co-partisan employers (Colonnelli2022). Non-monetary rewards, such as ego rents or prosocial benefits, may also play a role (gulzar2024good). Finally, to engage in politics, the individual incurs entry costs $c^{P}$ (e.g., campaign costs or membership registration fees).
@@
Second, the effects increase with the expected benefits of political engagement during unemployment, $(1-\lambda)b^{P}$, weighted by the time not already allocated to politics, $1-h^{P}$. Intuitively, this term reflects the notion that, when remaining unemployed, individuals can reallocate time that would otherwise be spent in market work towards political activities, thereby gaining additional political benefits. To shed light on one of the monetary returns from engaging with politics, we will investigate whether the treatment effects are larger in localities where councilors earn higher salaries. [^15] In addition to monetary returns, political involvement may also provide policy payoffs. In citizen-candidate models (osborne1996; besley1997), political candidates represent particular policies and care about winning per se. As argued by hall2019 and thomsen2017, the policy returns may be especially relevant for ideologically extreme candidates. Thus, we will examine potential heterogeneity in the effects on memberships and candidacies in parties of different ideological positions. Moreover, we will test the role of time resources by exploiting discontinuities in eligibility for UI benefits that increase laid-off workers’ incentives to remain unemployed.
@@
The framework also provides predictions about the potential implications of layoffs for the quality of individuals who select into politics. Given that the participation effects of layoffs are predicted to be larger for individuals with higher income losses and higher expected benefits from political engagement, the consequences for selection depend on how these factors are related to individual competence. High-ability individuals – such as those with higher education – tend to earn higher wages in the private sector. Thus, a layoff may represent a larger shock to their opportunity costs, leading to a more pronounced increase in their incentives to enter politics. In addition, the expected benefits of political engagement may vary with individuals’ competence. For instance, the returns to running for office rise with the probability of winning, which may be higher for more capable individuals. Moreover, the expected likelihood of gaining access to a public-sector job through political connections may depend on ability, either negatively if political support acts as a substitute for individual competence (colonnelli2020) or positively if politicians use their discretionary power to appoint more competent individuals (brollo2017; bazzi2023).
@@ 3.2 Labor market data
We obtain matched employer-employee data covering the universe of formal employees between 2004 and 2018 from the Relação Anual de Informações Sociais (RAIS), an administrative database collected annually by the Ministry of Labor and Employment. [^16] In 2015, formally employed individuals accounted for approximately 55% of the Brazilian labor force (derenoncourt2021). The database was created in 1975 for statistical and administrative purposes and contains detailed information such as hiring and separation date, average monthly earnings, contracted hours, occupation, firms’ location and industry, and workers’ socio-demographic characteristics. For our analysis, we make use of the identified dataset, containing individuals’ names, taxpayer registry numbers (CPF), as well as firms’ tax identification numbers (Cadastro Nacional da Pessoa Jurídica, CNPJ). This enables us to identify individuals’ employment histories over time and across firms.
@@ 3.3 Merging political and labor market data
We merge the information on political candidates with the employment data using an individual’s taxpayer registry number (CPF). As for party membership, we merge this information to the RAIS data based on each individual’s full name. To minimize measurement error when exactly matching on names, we follow Britto2022a; Britto2022 and Bhalotra2021 and restrict our sample to individuals who have a unique name in Brazil. Individuals with unique names cover roughly half of the Brazilian population (50.2% in our sample), given that individuals commonly have several surnames, including at least one from both their father and mother. [^17] To identify unique names, we build a dataset of names in Brazil. We collect the names of all workers who appear in the RAIS data in any year between 2004 and 2018, and supplement this list with information on the universe of beneficiaries in the social programs Bolsa Família and Benefício de Prestação Continuada between 2013 and 2021 and on all party members between 2000 and 2020. The resulting list covers about 90% of the Brazilian adult population. Not observing the names of a small part of the population will introduce some measurement error in our merged outcome variables, which is, however, not expected to be related to the likelihood of being laid off.
@@ 4.1 Empirical strategy
We aim to estimate the impact of adverse economic shocks on individuals’ likelihood of entering politics. To do so, we study individuals who are laid off from formal employment in the Brazilian labor market. Brazil’s private labor market exhibits very high job turnover (Gonzaga2003), with layoffs without just cause accounting for about three-quarters of all separations (and voluntary quits by workers accounting for most of the remaining quarter). Among all workers who appear in RAIS between 2004 and 2018, 66.6% experienced at least one layoff during this period. [^18] The yearly layoff rate (among those with a formal job in a given year) is 19.1%. Political entrants are not shielded from these shocks: among all individuals who newly became party members (ran for local councilor at least once) between 2000 and 2020, 36.8% (41.5%) experienced at least one layoff. [^19] Note that only 55.6% (70.3%) of all new party members (councilor candidates) ever appeared in RAIS with a formal contract over the 2004-2018 period. Thus, layoffs are common labor market shocks faced by individuals, including those who engage in politics.
@@
The main difficulty in estimating the effect of layoffs on individuals’ entry into politics is the fact that layoffs are not random. On the one hand, individual shocks could simultaneously influence both political participation and the likelihood of losing a job, leading to a spurious relationship. On the other hand, individuals’ political behavior could also affect their probability of being dismissed. Colonnelli2022 show, for instance, that political discrimination affects labor market outcomes in Brazil. To address these identification challenges, we exploit the timing of firm-level mass layoffs as a source of plausibly exogenous variation in individual-level job loss using a difference-in-difference design in a matched sample of workers. The timing of mass layoffs is arguably orthogonal to individual workers’ political behavior and other confounding worker-level shocks, and has been widely used to estimate the effects of job loss on various outcomes. In particular, our empirical strategy closely follows recent papers that study the effect of dismissals in Brazil (Bhalotra2021; Britto2022; Britto2022a; Amorim2023). This subsection describes how we construct the estimation sample and then presents the estimation model.
@@
Sample selection. We restrict our sample to workers aged 25 to 50 with a full-time (i.e., at least 30 hours per week), open-ended, private-sector job of at least six months tenure. Among those, the treatment group comprises all workers displaced without just cause from a mass layoff firm between 2009 and 2011.202020We exclude individuals who are recalled by their initial employer within five years after the layoff. In cases where individuals experienced multiple mass layoffs between 2009 and 2011, we keep only the first one. This treatment period allows us to study long-run effects over three election cycles (12 years) after the layoff, and also to check for pre-trends over two election cycles (8 years) before the layoff. In our baseline specification, we define a mass layoff firm as a firm with at least 30 workers dismissing at least 30% of its workforce in a given year, and later we show that our results are robust to various thresholds of firm size and layoff share. [^21] We always drop firms from the mass layoff sample if they reallocate under a new identifier, specifically if at least 30% of their dismissed workers transition to the same new firm identifier in the following year.
@@
We then match each treated individual with a control worker who is not dismissed in the same year and who works in a non-mass layoff firm. [^22] Note that even though control workers are not laid off in the matching year, they may be laid off in later years. Given that we estimate effects for up to 12 years after the layoff, restricting control workers to those who were continuously employed throughout the whole observation period would result in a small and highly selected pool of control workers. Defining control workers without conditioning on future employment follows a large share of papers in the job loss literature (e.g., Britto2022; bertheau2023; schmieder2023). In addition, we do not impose any restrictions on potential layoffs prior to the matching year, other than the minimum six-month tenure requirement. Given the high turnover rates in Brazil, stricter tenure requirements would yield a substantially selected sample. In Appendix Figure B.1, we show that treated and matched control individuals were not differently exposed to mass layoffs before the treatment period. To obtain a comparable control group, we perform an exact matching on gender, education (5 categories), age (deciles), monthly earnings (deciles), tenure (deciles), firm size (deciles), and state (27 categories). [^23] We allow control individuals to be matched to more than one treated individual. In cases where more than one control individual was matched to one treated individual, we randomly selected one control individual. For 79.5% of the workers in the initial treatment pool, we are able to find a comparable control worker. We end up with an estimation sample that consists of nearly one million treated and matched control workers.
@@
Table 1 presents summary statistics for treated and control individuals and shows that both groups are similar in socio-demographic, job, and firm characteristics. For all considered variables, the standardized mean difference between the two groups is well below the threshold of 0.20 recommended by imbens2015. Importantly, treated and control units are also balanced in terms of pre-treatment political outcome variables which were not part of the matching process. In both groups, approximately 7% of individuals were members of a political party in 2008. Nearly 2% newly registered as a member in the election cycle between 2005 and 2008, and 0.1% ran for local councilor in the 2008 election. [^24] Appendix Table B.1 also shows that there are no differences in the ideology, local incumbency, and union affiliation of the parties that treated and control observations affiliated with or ran for before the layoff. The validity of our DiD design does not require that treated and control individuals are similar in levels, but covariate balance makes it more plausible that both groups exhibit similar trends in counterfactual outcomes.
@@
which summarizes the average treatment effects over all periods in the coefficient $\delta_{DiD}$. In both models, we include municipality $\times$ two-digit industry $\times$ cycle fixed effects, $\mu_{m(i)k(i)c}$, which flexibly control for local and industry shocks that may simultaneously cause mass layoffs and changes in political behavior. Importantly, the municipality $\times$ cycle fixed effects also hold constant the local political environment and allow us to separate individual-level supply-side effects from potential effects of mass layoffs on the demand for politicians among the local electorate. [^25] Note that the estimation models (1) and (2) rely on a single treatment period, as all treated workers were laid off in the 2012 cycle. Thus, our main specification is not susceptible to weighting issues that arise in staggered difference-in-differences models with heterogeneous treatment effects (dechaisemartin2020; goodmanbacon2021; sun2021).
@@
where $i$, $t$, and $t^{*}$ refer to workers, calendar years, and layoff years, respectively. [^27] For control workers, we assign the layoff year of their matched treated worker. $\delta_{k}$ capture the dynamic treatment effects for each event year after the layoff $k$. The model includes calendar year $\times$ layoff year fixed effects, $\lambda_{tt^{*}}$, to ensure that the identifying variation only comes from comparing treated to control individuals who have been matched in the same year. [^28] This stacking approach avoids the “forbidden comparisons” between late and early-treated individuals that can cause negative weights under heterogeneous treatment effects (cengiz2019; dube2023).